### v1

@@ Abstract
Phase-separated manganites provide a unique platform to study the dynamics of competing electronic and structural orders in correlated systems. In La0.275Pr0.35Ca0.375MnO3 (LPCMO), we use temperature-cycling Raman spectroscopy to uncover a previously unidentified regime of structural irreversibility, emerging from the interplay between lattice distortions and phase competition across the phase-separation and charge–orbital ordering temperatures. This irreversible behavior encodes a thermal-memory effect reflecting the system’s history-dependent energy landscape. Correlated magnetic and transport responses confirm the coupling between lattice and electronic degrees of freedom, revealing a new form of nonequilibrium phase dynamics in mixed-valence oxides. These results advance the understanding of metastability and memory phenomena in strongly correlated materials, opening pathways toward adaptive and neuromorphic functionalities in quantum materials.
Introduction
@@
Mixed-valence manganites exemplify the complexity inherent to strongly correlated oxides, wherein the interaction among spin, charge, orbital, and lattice degrees of freedom gives rise to emergent properties with tunable functionalities. The phase-separation (PS) regime intrinsic to these materials, characterized by the coexistence of ferromagnetic metallic (FMM) and antiferromagnetic insulating with charge and orbital ordering (AFMI-COO) regions, imparts a notable sensitivity to external perturbations. Modest magnetic, electric, or thermal stimuli can dynamically reconfigure the PS landscape, producing resistive-switching and memory effects that have been widely explored for neuromorphic and adaptive electronics [Sacanell2004, Sacanell2018, Levy2002, Quintero2007, Gomide2025, Ivan2024, Schulman2024, Paasonen2024, Jaman2025, Zhang2020, Lahteenlahti2021, li2022, Zhou2015, Milloch2024].
@@
Among the extensively studied manganites, La5/8-yPryCa3/8MnO3 (LPCMO) serves as a prototypical system to investigate PS phenomena. At intermediate Pr concentrations, LPCMO exhibits sub-micrometric coexistence of FMM and AFMI-COO phases over a wide temperature range, giving rise to a dynamic-like phase separation (DPS) regime between approximately 30 and 100 K [Sarma2004, Uehara, Ghivelder2005, Garcia2011, Dhakal2007, Diego24]. In this regime, unpinned interfacial walls fluctuate and respond readily to magnetic and electric fields [Uehara, Dhakal2007, Merten, Ghivelder2005], mechanical strain [Podzorov2001, Ahn2004, Ogimoto2005, Ward2009, Kim2021], pressure [Laukhin1997, Baldini2012], light [Deng2024, Lin2018], and the compound’s temperature history [kundhikanjana2015, Sarma2004, Sacanell2004, Ward2011, Sacanell2018].
@@
Beyond contrasting electronic and magnetic behavior, the two coexisting phases also differ structurally. While the FMM regions maintain the compound’s room-temperature orthorhombic Pnma symmetry, the AFMI-COO matrix exhibits a monoclinic P21/m structure, originating from cooperative Jahn-Teller (JT) distortions of Mn3+ ions and associated charge and orbital ordering [Collado2003, Radaelli1997, Goff2004, Garcia2011]. The contrast between long- and short-range JT distortions provides a direct structural fingerprint of the PS evolution. Raman spectroscopy, noted for its exceptional sensitivity to lattice distortions and orbital-phonon coupling, is particularly adept at probing these effects in real time [Granado1998, Granado2000, Kim, Merten, Dediu2000, Abrashev1999, Adams2000, Moshnyaga2014, Granado2001].
@@
Laser-based approaches extend the control landscape of correlated systems by directly engaging the fundamental lattice, spin, and orbital degrees of freedom that define their collective behavior. Ultrafast optical excitation can coherently drive phonons, modulate exchange interactions, and reshape electronic bandwidths on femtosecond timescales. Unlike electrical or magnetic stimuli, which act non-locally through current flow or field gradients, laser perturbations are intrinsically local and contactless, enabling precise spatial control with negligible dissipation. This selectivity grants access to hidden metastable configurations and non-equilibrium pathways unattainable under static driving, leading to light-induced resistive and magnetic switching [Deng2024, Lin2018]. Such all-optical schemes offer a powerful route toward quantum and neuromorphic devices, where coherence preservation, minimal Joule heating, and nanoscale phase engineering are critical for scalable operation [Kalcheim2020].
Here, we employ temperature-cycling Raman spectroscopy (Figure 1 a) to uncover a previously unreported regime of structural irreversibility associated with the reconfiguration of the DPS network in LPCMO. By applying cumulative, pulsed-like temperature variations between low base, TB, and target, TT, temperatures, depicted in figure 1 b, we observe metastable lattice states imprinted in the intensity and width of the JT stretching modes. Figure 1 c displays the “as-cooled” spectrum, collected before starting the thermal cycling procedure with laser power. Figures 1 d and e reveal that, measured at TB after successive thermal cycles between TB and TT = 102(4) and 124(5) K, both asymmetric (AS) and symmetric (S) JT distortion modes gain intensity relative to the others. For cycles with higher TT, the intensities of these modes decrease continuously up to TT = 188(5) K, vide figures 1 f - h. For this and higher values of TT (not shown), the Raman spectra after cycling are very similar to the “as-cooled” one. These irreversible changes occur only within a specific temperature window bounded by the DPS critical (TC) and charge-orbital ordering (TCOO) temperatures and vanish for T TCOO, revealing a distinct thermal-memory effect in the lattice degrees of freedom.
@@
Complementary magnetic and transport measurements corroborate that these structural modifications coincide with the reorganization of the PS network. Altogether, these findings demonstrate that targeted thermal and optical protocols can be used to manipulate the balance between coexisting FMM and AFMI-COO phases at will. This establishes a pathway for laser-assisted phase engineering in correlated oxides, bridging fundamental phase-transition physics with emergent opportunities in optically controlled, low-power quantum and neuromorphic devices.
Results
@@
Figure 2 presents the standard thermal (cold finger) evolution of the LPCMO Raman spectrum. Figure 2 a shows spectra collected at T = 38, 121, 221 and 321 K. The position of the main rotational ( 251 cm-1), bending ( 440 cm-1), asymmetric, AS-JT ( 485 cm-1) and symmetric, S-JT ( 615 cm-1), distortion modes [Carron, Granado2000, Kim, Iliev1998, Merten, Iliev2003, Granado1998, Liarokapis1999, Merten2019, Amelitchev2001] are indicated by black dashed vertical lines. The red dashed vertical line indicates the presence of a spurious peak, possibly due to local oxygen vacancies. Figure 2 b displays an intensity map of the Raman spectra, with indications of the transition temperatures of the sample represented by dashed horizontal lines. These spectra show the AS-JT and JT modes intensities increasing significantly around 100 K, which is accompanied by the appearance of extra rotational and bending modes around 220, 330 and 515 cm-1 [Kim, Merten], respectively. Figures 2 c and 2 d display the full widths at half maximum (FWHM) and intensities of the AS-JT and S-JT, respectively. The intensities were normalized by the rotational mode present at 250 cm-1, which was taken as a reference. These modes broaden and become stronger as the system evolves from the DPS to the AFMI-COO regimes. The maximum intensities of the JT modes occur around TMI. Above TCOO, these modes weaken substantially.
@@
Figure 3 shows further details of the evolution of the AS- and S-JT modes as a function of the thermal cycling process. Figure 3 a shows a color map of the Raman intensities, measured at T K, as a function of the target temperature TT. The similarity of this map with the one shown in figure 2 b is remarkable, despite the fact that these are completely different experimental protocols. The JT modes start to gain intensity around TT = TC, with maximum intensity close to TT = TMI, and then lose intensity as TT approaches TCOO ( 205 K). From TT = TCOO onward, the mode intensities return to their original values, i.e., they are reversed or reset. Figure 3 b shows the FWHM of the JT modes as a function of TT. As TT increases, the AS-JT mode undergoes a broadening followed by a sharpening, while the width of the S-JT mode shows less significant variation. This, in conjunction with the normalized mode intensities in figure 3 c, reveals that the thermal cycling process induces an enhancement, between TC TT TCOO, with a maximum around TMI, followed by a weakening of COO phase at TB. That is, the strength of the COO phase achieved at TT is retained when the system cools back to TB.
@@
The normalized intensities of both modes, in both protocols, figures 2 d and 3 c, show a quantitative correspondence for TT T TC. However, for TC TT T TCOO, such an agreement is not observed (apart from TT T 100 K in the AS-JT mode). Since what is being shown is the relative intensity between two distinct vibrational modes, this can be interpreted as a partial loss of information during the cooling process. The state reached at TB after thermal cycling for temperatures between TC TT TCOO does not fully recover the state to which it was taken due to heating. Hence, during the cooling process, and due to the high dynamism of the phases throughout the temperature range considered here [Ghivelder2005, Dhakal2007, Lee2002], a new rearrangement of the phases occurs. However, this is not sufficient to completely ”erase” the information stored by the material due to the thermal process to which it has been exposed.
@@
It is remarkable how, when performing the thermal cycling process at temperatures TC and TCOO, the system has the capacity to partially recover and reset the state it reached at the different target temperatures. With the process of cyclic temperature variation, characteristics of the vibrational modes found in a conventional heating process are reached and stored at the base temperature of the system. This capacity to store and erase seems to be closely related to the material’s well-known transitions [Ghivelder2005, Dhakal2007, Lee2002, Kiryukhin2000, Diego24]. Performing cumulative thermal cycles that transition from its DPS to its mostly AFMI-COO regime, TC, produces an intensity gain in its JT distortion modes that is similar to the gain observed when the compound is heated conventionally. Cyclically varying the temperature to TT TCOO has the effect of erasing this intensity gain and reverting the system to its initial state.
@@
In order to quantify these intensity changes, figures 3 d and e show the normalized difference, , between the initial and final intensity values, , considering each individual cycle (figure 3 d) with , and the initial ”as-cooled” intensity values (figure 3 e). The individual contribution to the modes intensities of each thermal cycle reaches its maximum between TC TT TMI. The intensity relation is reversed for TMI TT TCOO, i.e., starts to diminish in relation to Ii. For TT T TCOO, within the fitting error, no further changes are observed due to thermal cycling. In figure 2 e , it can be seen how, with respect to the initial ”as-cooled” JT mode intensity values, an almost two-fold gain is observed between TC and TMI. This gain diminishes as the target temperature approaches TCOO, and finally, it is reset after this transition, returning to values close to the initial ones.
@@
Complementary magnetization and resistivity measurements as a function of analogous thermal cycles were performed to correlate local Raman results with transport and magnetic macroscopic properties. Panel 4 a shows the full ”Cooling”/”Warming” magnetization curves as a function of temperature (M - T), collected by cooling from 300 to 10 K and heating from 10 to 300 K, together with different M - T curves between TB and the values of interest of TT. All these measurements were obtained with no applied DC-field. In the full M - T curve, it is possible to notice inflections at 205 K in heating and cooling protocols (LPCMO’s COO transition temperature, TCOO), and at 91 K and 63 K while warming and cooling, respectively. The latter corresponds to T and T, respectively. These transition temperatures were determined from the maximum and minimum points of the first derivative of the moment with respect to temperature, as shown in figure S8 a in the Supplementary Materials S.2.
@@
From the initial magnetization, denoted as Mi, and the final magnetization, denoted as Mf, at temperature TB preceding and following each thermal cycle, three distinct regimes can be observed. For TT 80 K (TT 70 K in 4 a), the final magnetization is approximately equal to the initial one, with this difference decreasing as TT approaches 80 K. For TT 90 K, the final magnetization begins to be smaller than the initial one until both values become approximately equal at TT = 150 K. This persists until TT 190 K, from where the difference between Mf and Mi begins to increase again until TT 210 K, when a considerable jump in Mf compared to Mi is observed. As TT increases, the difference between the final and initial magnetization approaches zero again.
@@
In panel 4 b resistivity curves as a function of temperature ( - T) were collected by cooling and warming between 300 and 10 K, together with - T curves between TB and values of interest of TT are shown, analogously to panel 4 a. The resistivity measurements show an insulator-metal transition at approximately 120 K during cooling (TIM). During heating, a metal-insulator transition occurs at around 153 K (TMI). The positions of these were determined by the sign of the first derivative of the resistivity as a function of temperature and can be seen in figure S8 b in the Supplementary Materials S.2.
@@
By comparing the thermal dependence of magnetic (4 a) and electrical (4 b) properties, it is possible to see that the final temperature at which the decrease (or increase) in final magnetization in relation to the initial magnetization after each cycle corresponds to the increase (or decrease) in final resistivity in relation to the initial resistivity. The panel 4 c illustrates this comparable and complementary behavior with M and measured at TB before (Mi and ) and after (Mf and ) each thermal cycle. The aforementioned correlation between M and is immediately observed. Furthermore, both cases show a dependence on the final state after each cycle with the transition temperatures TC, TMI and TCOO, as in the intensity and width of the JT distortion modes.
@@
Panel 4 d shows, analogously to panel 2 d, the normalized change ( in M and at TB with respect to the individual cycles (left, ) and with respect to the initial values after cooling to TB and before starting the thermal cycles (right, ). These figures, especially the one on the left of panel 4 d, show once again how TC and TCOO are the main factors responsible for the behavior presented thus far. In the right figure of panel 4 d it is possible to see a two-fold resistivity gain between TC and TCOO, when compared to the initial value. However, sample magnetization does not show quantitative agreement with these changes, producing only an approximate decrease in half in the same TT interval. This is possibly related to the AFM character of the insulating phase, which suppresses a two-fold decrease in magnetization.

### v2

@@ Abstract
Phase-separated manganites provide a unique platform to study the dynamics of competing electronic and structural orders in correlated systems. In La0.275Pr0.35Ca0.375MnO3 (LPCMO), we use temperature-cycling Raman spectroscopy to uncover a previously unidentified regime of structural irreversibility, emerging from the interplay between lattice distortions and phase competition across the phase-separation and charge–orbital ordering temperatures. This irreversible behavior encodes a thermal-memory effect reflecting the system’s history-dependent energy landscape. Correlated magnetic and transport responses confirm the coupling between lattice and electronic degrees of freedom, revealing a new form of nonequilibrium phase dynamics in mixed-valence oxides. These results advance the understanding of metastability and memory phenomena in strongly correlated materials.
@@
Mixed-valence manganites exemplify the complexity inherent to strongly correlated oxides, wherein the interaction among spin, charge, orbital, and lattice degrees of freedom gives rise to emergent properties with tunable functionalities. The phase-separation (PS) regime intrinsic to these materials, characterized by the coexistence of ferromagnetic metallic (FMM) and antiferromagnetic insulating with charge and orbital ordering (AFMI-COO) regions, imparts a notable sensitivity to external perturbations. Modest magnetic, electric, or thermal stimuli can dynamically reconfigure the PS landscape, producing resistive-switching and memory effects that have been widely explored for neuromorphic and adaptive electronics [47, 48, 32, 45, 17, 8, 50, 43, 23, 55, 29, 34, 56, 40].
@@
Among the extensively studied manganites, La5/8-yPryCa3/8MnO3 (LPCMO) serves as a prototypical system to investigate PS phenomena. At intermediate Pr concentrations, LPCMO exhibits sub-micrometric coexistence of FMM and AFMI-COO phases over a wide temperature range, giving rise to a dynamic-like phase separation (DPS) regime between approximately 30 and 100 K [49, 52, 15, 14, 12, 7]. In this regime, unpinned interfacial walls fluctuate and respond readily to magnetic and electric fields [52, 12, 39, 15], mechanical strain [44, 3, 42, 53, 26], pressure [30, 5], light [11, 36], and the compound’s thermal history [28, 49, 47, 54, 48].
@@
Beyond contrasting electronic and magnetic behavior, the two coexisting phases also differ structurally. While the FMM regions maintain the compound’s room-temperature orthorhombic Pnma symmetry, the AFMI-COO matrix exhibits a monoclinic P21/m structure, originating from cooperative Jahn-Teller (JT) distortions of Mn3+ ions and associated charge and orbital ordering [9, 46, 16, 14]. The contrast between long- and short-range JT distortions provides a direct structural fingerprint of the PS evolution. Raman spectroscopy, noted for its exceptional sensitivity to lattice distortions and orbital-phonon coupling, is particularly adept at probing these effects in real time [19, 20, 25, 39, 10, 1, 2, 41, 18].
@@
Laser-based approaches extend the control landscape of correlated systems by directly engaging the fundamental lattice, spin, and orbital degrees of freedom that define their collective behavior. Ultrafast optical excitation can coherently drive phonons, modulate exchange interactions, and reshape electronic bandwidths on femtosecond timescales. Unlike electrical or magnetic stimuli, which act non-locally through current flow or field gradients, laser perturbations are intrinsically local and contactless, enabling precise spatial control with negligible dissipation. This selectivity grants access to hidden metastable configurations and non-equilibrium pathways unattainable under static driving, leading to light-induced resistive and magnetic switching [11, 36]. Such all-optical schemes offer a powerful route toward quantum and neuromorphic devices, where coherence preservation, minimal Joule heating, and nanoscale phase engineering are critical for scalable operation [24].
@@
Here, we employ temperature-cycling Raman spectroscopy to uncover a previously unreported regime of structural irreversibility associated with the reconfiguration of the DPS network in LPCMO. By applying cumulative, pulsed-like temperature variations between low base, TB, and target, TT, temperatures, we observe metastable lattice states imprinted in the intensity and width of the JT stretching modes. Complementary magnetic and transport measurements corroborate that these structural modifications coincide with the reorganization of the PS network. Altogether, these findings demonstrate that targeted thermal and optical protocols can be used to manipulate the balance between coexisting FMM and AFMI-COO phases at will. This establishes a pathway for laser-assisted phase engineering in correlated oxides.
Polycrystalline LPCMO () samples were prepared using a standard solid-state reaction method, as described by [9, 33, 7]. Its orthorhombic Pnma (a c b) single phase character at room temperature was confirmed by synchrotron x-ray powder diffraction performed in the PAINEIRA beamline in CNPEM-SIRIUS [13, 9]. The Rietveld refinement performed with the GSAS-II software [51] and the refined crystallographic parameters can be found in figure S7 and table S2 in Supplementary Materials S.
Raman scattering experiments were performed in a quasi-backscattering geometry using a 532 nm laser line with a spot size at the sample of 50 m, as depicted in figure 1. A Jobin-Yvon T64000 spectrometer with a single 600 mm-1 grating and a set of Bragg notch filters was used. A LN2-cooled multichannel charge-coupled device (Symphony II, Horiba) collected the scattered data. The sample was glued on the cold finger of a closed-cycle He cryostat with a minimal temperature of K. Spectra as a function of the cold-finger nominal temperature were collected while heating to investigate the thermal evolution of the sample’s vibrational modes and to serve as a control parameter for the measurements performed as a function of laser power. This protocol will be referred to as ”standard procedure/heating”. In the measurements as a function of laser power, the sample was first cooled and maintained at the cryostat minimal temperature. Then, two different protocols were employed: 1) a continuous increase in power, with spectra collection using 10 mW to 150 mW (with 10 mW steps), and 2) cycling the power, where spectra were collected at 10 mW (base power) before and after it was increased to target powers ranging from 20 mW to 150 mW, illustrated in figure 1. The latter method of temperature variation allows for almost instantaneous heating and subsequent cooling of the analyzed region, and will be referred as ”cyclic procedure/heating”. The comparison between the spectra collected in protocol 1) and those collected in the standard procedure yielded the temperature of the analyzed spot as a function of power, as detailed in the Supplementary Materials S. The thermal cycles were conducted between TB = 38(4) K and TT = 59(4), 81(4), …, 317 (7) K; see table S1 in the Supplementary Materials S. In both experiments, the laser spot probed the same position on the sample to enhance control over the localized heating effects induced by the laser, and the collection time was proportionally decreased as the laser power increased. Different spots and pieces of the sample were measured with the same protocol.
Magnetic and electric characterizations were performed using a MPMS 3 and PPMS Model 6000 (Quantum Design), respectivelly. Magnetization as a function of temperature, M-T, was measured in the VSM mode with an applied field of 50 Oe. Extractions were performed between 300 K and 2 K, cooling and warming at a rate of 2 K/min. The sample resistance was measured using a standard two-probe technique with spacing between analyzer probes of 0.7 mm and a compliance current of 0.1 A. Resistance as a function of temperature, R-T, was measured between 300 K and 10 K, cooling and heating at a rate of 2 K/min, respectively. Magnetization and resistance as a function of the cyclic procedure between a base, TB = 30 K, and target, TT, 50, 60, 70, … 250 K, temperatures were measured after cooling to 30 K. Warming and cooling M-T and R-T curves between TB and TT were collected with a cooling/warming rate of 5 K/min.
@@
Figure 2 presents the standard thermal (cold finger) evolution of the LPCMO Raman spectrum. Figure 2 (a) shows spectra collected at T = 38, 121, 221 and 321 K. The position of the main rotational ( 251 cm-1), bending ( 440 cm-1), asymmetric, AS-JT ( 485 cm-1) and symmetric, S-JT ( 615 cm-1), distortion modes [37, 20, 25, 21, 39, 22, 19, 35, 38, 4] are indicated by black dashed vertical lines. The red dashed vertical line indicates the presence of a spurious peak, possibly due to local oxygen vacancies. Figure 2 (b) displays an intensity map of the Raman spectra, with indications of the transition temperatures of the sample represented by dashed horizontal lines. These spectra show the AS-JT and JT modes intensities increasing significantly around 100 K, which is accompanied by the appearance of extra rotational and bending modes around 220, 330 and 515 cm-1 [25, 39], respectively. Figures 2 (c) and 2 (d) display the full widths at half maximum (FWHM) and intensities of the AS-JT and S-JT, respectively. The intensities were normalized by the rotational mode present at 250 cm-1, which was taken as a reference. These modes broaden and become stronger as the system evolves from the DPS to the AFMI-COO regimes (TC). The maximum intensities of the JT modes occur around the metal-insulator transition temperature, TMI. Above TCOO, the charge and orbital ordering temperature, these modes weaken substantially.
Figure 3 (a) displays the “as-cooled” spectrum, collected before starting the thermal cycling procedure with laser power, in black, as well as spectra collected at TB after after successive thermal cycles between TB and TT = 102(4), 124(5), 167 (5) and 188 (5) K. Is revealed that both asymmetric (AS) and symmetric (S) JT distortion modes gain intensity relative to the others until 145 (5) K. For cycles with higher TT, the intensities of these modes decrease continuously up to TT = 188(5) K. For this and higher values of TT (not shown), the Raman spectra after cycling are very similar to the “as-cooled” one. These irreversible changes occur only within a specific temperature window bounded by the DPS critical (TC) and charge-orbital ordering (TCOO) temperatures and vanish for T TCOO, revealing a distinct thermal-memory effect in the lattice degree of freedom.
@@
Figure 3 (b) shows a color map of the Raman intensities at as a function of the laser-induced local target temperature . The similarity of this map with the one shown in figure 2 (b) is remarkable, despite the fact that these are completely different experimental protocols. The JT modes start to gain intensity around TT = TC, with maximum intensity close to TT = TMI, and then lose intensity as TT approaches TCOO ( 205 K). From TT = TCOO onward, the mode intensities return to their original values, i.e., they are reversed or reset. Figure 3 (c) shows the FWHM of the JT modes as a function of TT. As TT increases, the AS-JT mode undergoes a broadening followed by a sharpening, while the width of the S-JT mode shows less significant variation. This, in conjunction with the normalized mode intensities in figure 3 (d), reveals that the thermal cycling process induces an enhancement, between TC TT TCOO, with a maximum around TMI, followed by a weakening of COO phase at TB. That is, the strength of the COO phase achieved at TT is retained when the system cools back to TB.
@@
The normalized intensities of both modes, in both protocols, figures 2 (d) and 3 (d), show a quantitative correspondence for TT T TC. However, for TC TT T TCOO, such an agreement is not observed (apart from TT T 100 K in the AS-JT mode). Since what is being shown is the relative intensity between two distinct vibrational modes, this can be interpreted as a partial loss of information during the cooling process. The state reached at TB after thermal cycling for temperatures between TC TT TCOO does not fully recover the state to which it was taken due to heating. Hence, during the cooling process, and due to the high dynamism of the phases throughout the temperature range considered here [15, 12, 31], a new rearrangement of the phases occurs. However, this is not sufficient to completely ”erase” the information stored by the material due to the thermal process to which it has been exposed.
@@
It is remarkable how, when performing the thermal cycling process at temperatures TC and TCOO, the system has the capacity to partially recover and reset the state it reached at the different target temperatures. With the process of cyclic temperature variation, characteristics of the vibrational modes found in a conventional heating process are reached and stored at the base temperature of the system. This capacity to store and erase seems to be closely related to the material’s well-known transitions [15, 12, 31, 27, 7]. Performing cumulative thermal cycles that transition from its DPS to its mostly AFMI-COO regime, TC, produces an intensity gain in its JT distortion modes that is similar to the gain observed when the compound is heated conventionally. Cyclically varying the temperature to TT TCOO has the effect of erasing this intensity gain and reverting the system to its initial state.
@@
In order to quantify these intensity changes, figures 3 (e) and (f) show the normalized difference, , between the initial and final intensity values, , considering each individual cycle (figure 3 (e)) with , and the initial ”as-cooled” intensity values (figure 3 (f)). The individual contribution to the modes intensities of each thermal cycle reaches its maximum between TC TT TMI. The intensity relation is reversed for TMI TT TCOO, i.e., starts to diminish in relation to Ii. For TT T TCOO, within the fitting error, no further changes are observed due to thermal cycling. In figure 2 (f) , it can be seen how, with respect to the initial ”as-cooled” JT mode intensity values, an almost two-fold gain is observed between TC and TMI. This gain diminishes as the target temperature approaches TCOO, and finally, it is reset after this transition, returning to values close to the initial ones.
@@
Complementary magnetization and resistivity measurements as a function of analogous thermal cycles were performed to correlate local Raman results with transport and magnetic macroscopic properties. Figures 4 (a) and (b) shows the full ”Cooling”/”Warming” magnetization curves as a function of temperature (M - T), collected by cooling from 300 to 10 K and heating from 10 to 300 K, together with different M - T curves between TB and the values of interest of TT, respectively. In the full M - T curve, it is possible to notice inflections at 205 K in heating and cooling protocols (LPCMO’s COO transition temperature, TCOO, see figures 4 (a) inset), and at 91 K and 63 K while warming and cooling, respectively. The latter corresponds to T and T, respectively. These transition temperatures were determined from the maximum and minimum points of the first derivative of the moment with respect to temperature, as shown in figure S8 (a) in the Supplementary Materials S.2.
@@
From the initial magnetization, denoted as Mi, and the final magnetization, denoted as Mf, at temperature TB preceding and following each thermal cycle, distinct regimes can be observed in figure 4 (b). For TT 80 K (not shown), the final magnetization is approximately equal to the initial one, with this difference decreasing as TT approaches 80 K. For TT 90 K, the final magnetization begins to be smaller than the initial one until both values become approximately equal at TT = 150 K. This persists until TT 190 K, from where the difference between Mf and Mi begins to increase again until TT 210 K, when a considerable jump in Mf compared to Mi is observed. As TT increases, the difference between the final and initial magnetization approaches zero again.
@@
In panel 4 (c) and (d) resistivity curves as a function of temperature ( - T) were collected by cooling and warming between 300 and 10 K, together with - T curves between TB and values of interest of TT are shown, analogously to figures 4 (a) and (b). The resistivity measurements show an insulator-metal transition at approximately 120 K during cooling (TIM). During heating, a metal-insulator transition occurs at around 153 K (TMI). The positions of these were determined by the sign of the first derivative of the resistivity as a function of temperature and can be seen in figure S8 (b) in the Supplementary Materials S.2.
@@
By comparing the thermal dependence of magnetic (figures 4 (a) and (b)) and electrical (4 (c) and (d)) properties, it is possible to see that the final temperature at which the decrease (or increase) in final magnetization in relation to the initial magnetization after each cycle corresponds to the increase (or decrease) in final resistivity in relation to the initial resistivity. The figures 5 (a) and (b) illustrates this comparable and complementary behavior with M and measured at TB before (Mi and ) and after (Mf and ) each thermal cycle. The aforementioned correlation between M and is immediately observed. Furthermore, both cases show a dependence on the final state after each cycle with the transition temperatures TC, TMI and TCOO, as in the intensity and width of the JT distortion modes.
@@
Figures 5 (c) and (d) shows, analogously to figures 2 (e) and (f), the normalized change ( in M and at TB with respect to the individual cycles (figure 5 (c), ) and with respect to the initial values after cooling to TB and before starting the thermal cycles (figure 5 (d), ). These figures, especially figures 4 (a) and (b), show once again how TC and TCOO are the main factors responsible for the behavior presented thus far. In figure 5 (d) it is possible to see a two-fold resistivity gain between TC and TCOO, when compared to the initial value. However, sample magnetization does not show quantitative agreement with these changes, producing only an approximate quarter in the same TT interval. This is possibly related to the percolative nature of the FMM regions, which causes the non compatible changes in resistivity and magnetization.